Derive the following reduction formula :
Integral x^n e^x dx=x^n e^x-n integral x^n-1 e^x dx
1
Expert's answer
2020-05-13T19:54:54-0400

Given

"\int x^n e^xdx"

Let

"\begin{aligned}
u&=x^n \ \ \ \ \ \ \ &dv=&e^xdx\\
du&=nx^{n-1}dx \ \ \ \ \ \ \ &v=&e^x
\end{aligned}"

Then

"\begin{aligned}
\int x^n e^xdx&= uv-\int vdu\\
&=x^n e^x- n \int x^{n-1} e^xdx 
\end{aligned}"